OpenStudy (anonymous):

Can someone verify that the sign would be a negative or positive? (9m)^-2 = -81m is what I came up with.

OpenStudy (anonymous):

Hi Satellite :)

OpenStudy (anonymous):

take the reciprocal and square it, do not make it negative hello

OpenStudy (anonymous):

so it's 81m

OpenStudy (anonymous):

math is weird but not that weird if it is supposed to be negative you would see a big fat minus sign out front

OpenStudy (anonymous):

well, i am trying to get the hang of the signs ...doing my best lol

OpenStudy (anonymous):

\[b^{-n}=\frac{1}{b^n}\] so \[(9m)^{-2}=\frac{1}{(9m)^2}\]

OpenStudy (anonymous):

wow, way off. lol

OpenStudy (anonymous):

the minus sign in the exponent does not mean make it negative the minus sign in the exponent means take the reciprocal

OpenStudy (anonymous):

so for example \[3^{-2}=\frac{1}{9}\]

OpenStudy (anonymous):

so its 1/(9m)^2

OpenStudy (anonymous):

says simplify is that simplified?

OpenStudy (anonymous):

you could write \[\frac{1}{81m^2}\]if you like